# Exergetic Port-Hamiltonian Systems Modeling Language

### v1

@@ 5 Exergy
Every system is ultimately determined by a power-preserving interconnection of primitive subsystems, also called components, which represent elementary physical behaviors such as storage and reversible as well as irreversible exchange of energy. To achieve a thermodynamically consistent combination of reversible and irreversible dynamics, components are defined with respect to a fixed exergy reference environment that is discussed in this section.
@@ 5.1 Intuition
As briefly discussed in Section 1.4, thermodynamic models do not have all their physically-relevant degrees of freedom determined by their state variables, which leads to irreversibility. According to the first and second law of thermodynamics, internal energy, which is given by an imprecise phenomenological description, cannot be fully converted into mechanical or electromagnetic energy, which is precise, in the sense that the relevant degrees of freedom are determined by the state variables.
The irreversible degradation of ‘mechanically precise’ forms of energy into ‘thermodynamically imprecise’ forms of energy can be quantified using the concept of exergy. A hypothetical conversion device, whose mode of operation is only restricted by the first and second law of thermodynamics is used to determine how much mechanical (or electromagnetic) energy can be extracted from a system before it reaches thermodynamic equilibrium with a reference environment, whose intensive variables, such as temperature and pressure, are constant (in time and space), see Figure 9.
@@
Crucially, the ideal conversion device is not restricted to interacting with the system through its interface. Instead, the device operates in a counterfactual where all energy domains of both the system and the environment can be accessed. This means that the exergy content of a (hierarchically flattened) composite system can be obtained by summing the exergy stored in each of its storage components. So, in particular, the exergy content of a system does not depend on how its subsystems are interconnected. The exergy content of a component storing internal energy is assessed based on an ideal conversion device and the reference environment, while the exergy content of a component storing only mechanical or electromagnetic forms of energy is equal to its energy content.
As also remark that a similar concept of available energy could be defined intrinsically, meaning without relying on a reference environment. The reversible conversion device would then interact only with the system of interest until it reaches equilibrium with itself. However, the use of a reference environment with fixed intensive quantities has the important advantage that the exergy content of a system is given explicitly by its exergy storage function, which follows systematically from its energy storage function and the definition of the reference environment, as detailed in Section 6.1. In contrast, the intrinsic concept in general would lead to an optimization problem without a simple closed-form solution.
More information and references about exergy, its history, and its links to port-Hamiltonian as well as GENERIC systems can be found in Lohmayer et al. (2021).
@@ 5.2 Reference environment
A reference environment $(I_{R},\,\lambda_{R})$ is given by an interface $I_{R}=(N_{R},\,\tau_{R})$ such that each port $p\in I_{R}$ is a power port with an unique extensive quantity in $\mathbf{Q}_{\mathbb{R}}$ and an associated value $\lambda_{R}(p)\in\mathbb{R}^{*}$ of the energy-conjugate intensive quantity. ∎
@@
The minimal definition of a reference environment represents an isothermal heat bath. This amounts to an interface with a single entropy port and an associated reference temperature, henceforth denoted by ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$:
$N_{R}\>$ =$=\>\{\mathtt{s}\}$
$\tau_{R}(\mathtt{s})\>$ $=\>\left((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p}\right)$
@@
To determine the exergy content of pressurized fluids with respect to an isobaric atmosphere, the definition of the environment can be extended to $N_{R}=\{\mathtt{s},\,\mathtt{v}\}$ with
$\tau_{R}(\mathtt{v})\>$ $=\>\left((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p}\right)$
[…]
The minus sign in front of the reference pressure ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}}$ stems from the fact that more volume means less energy. To model for instance a mass diffusion process, a port $\mathtt{m}$ representing the mass of a chemical species with a fixed chemical potential ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\mu_{0}}$ should be included, i.e. let $N_{R}=\{\mathtt{s},\,\mathtt{v},\,\mathtt{m}\}$ with
[…]
A software implementation should include a comprehensive definition of a reference environment and allow the user to redefine the values of the intensive quantities.
@@ 6 Primitive systems
The lowest level of any hierarchical model specification is defined by primitive systems, also called components. There are four classes of components:
@@
    (or a transformation between different forms of describing an energy domain).
@@
    Environment components are used to model energy exchange with
@@
Components are in general defined with respect to a reference environment $(I_{R},\,\lambda_{R})$. Further, every component has an interface $I=(N,\,\tau)$ with port variables $(x,\,f,\,e)\in\mathcal{P}_{I}$. To define the exergy storage function of storage components and to state fundamental thermodynamic properties of reversible as well as irreversible components, we define the following notation: For any port $p\in I$, let
@@
For instance, if $p$ represents a thermal energy domain then $\lambda(p)={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$, whereas if $p$ represents a kinetic energy domain then $\lambda(p)=0$, since the environment does not have a kinetic energy domain. This can also be applied to the entire interface. E.g. for $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
@@
We also define a similar notation that references a specific environment port: For any port $p\in I$ and $p_{R}\in I_{R}$, let
@@
Again, given e.g. $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
@@ 6.1 Storage components
Storage components are primitive systems representing energy storage. Examples include elastic springs, movable masses, capacitors, inductors, and thermal capacities.
@@
The function $E$ simply yields the stored energy for any state $x\in\mathcal{X}_{I}$. Based on the reference environment $I_{R}$, the exergy content is determined as follows.
@@
Let $(I,\,E)$ be a storage component. Its exergy storage function $H:\mathcal{X}_{I}\rightarrow\mathbb{R}$ is given by
@@
$\forall\,p\in I\>:\>p\mathtt{.e}\>=\>\left.\partialderivative{H}{(p\mathtt{.x})}\right|_{x}$
In case the constant offset should matter for thermodynamic analysis, it can be added to the total exergy function once the equilibrium state of the system under study has been determined. Finally, the trajectory of the state variables is obtained by integrating the flow variables:
@@
Consider a storage component $(I_{\text{gas}},\,E_{\text{gas}})$ representing a gas-filled compartment whose entropy and volume can change. Its interface $I_{\text{gas}}$ is defined as in 2.18 and we assume knowledge of its energy storage function $E_{\text{gas}}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$. Writing $x=(s,\,v)$, the exergy storage function is then given by
$H_{\text{gas}}(s,\,v)\>=\>E_{\text{gas}}(s,\,v)\,-\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,s\,+\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}}\,v\,.$
With the temperature being defined as $\theta=\partialderivative{E_{\text{gas}}}{s}$ and the negative pressure being defined as $-\pi=\partialderivative{E_{\text{gas}}}{v}$, the effort variables are given by
$\begin{split}\mathtt{s.e}\>&=\>\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\\ \mathtt{v.e}\>&=\>-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}})\,.\end{split}$
Since the entropy rate $\dot{s}=\mathtt{s.f}$ is related to the heat flow rate $\Phi$ via $\dot{s}=\frac{1}{\theta}\,\Phi$, the exergetic power at the entropy port is $\langle\mathtt{s.e}\mid\mathtt{s.f}\rangle=\frac{\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}}{\theta}\,\Phi=\eta_{\text{Carnot}}\,\Phi$, where $\eta_{\text{Carnot}}$ is the efficiency of an ideal heat engine operating between a hot source with temperature $\theta$ and a cold sink with temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$. Considering a piston with the pressurized gas on one side and the atmosphere on the other, $\langle\mathtt{v.e}\mid\mathtt{v.f}\rangle=-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}})\,\dot{v}$ is the power that can be extracted from the pressurized gas when its volume grows at rate $\dot{v}$. ∎
Next, we define the semantics of a storage component with interface $I$ in terms of a relation $\mathcal{S}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$. This is used in Section 7, where we define the semantics of a composite system in terms of the semantics of its primitive subsystems and the semantics of its (hierarchy of nested) interconnection pattern(s).
@@ 6.2 Environment components
While the primary role of the reference environment is to enable the thermodynamically consistent combination of reversible and irreversible dynamics, it may also be represented by a component to model energy exchange with the environment. For instance, an environment component can represent an isothermal heat bath, which may absorb heat generated by irreversible components, or an isobaric atmosphere, which may exchange pressure-volume work with other components.
@@
The state variables keep track of the exchanged entropy, volume, etc. As for storage components, their rate of change is given by the flow variables:
@@
Since the exergy content of the environment is zero by definition, the effort variables are also zero:
@@
Consequently, the semantics of an environment component is assigned as follows:
@@
The semantics assigned to environment component $I\subseteq I_{R}$ is the relation $\mathcal{E}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$ that is given by the subbundle
@@ 6.3 Reversible components
In a generalized sense, reversible components represent gyrators, transformers, and/or kinematic constraints. Generalized gyrators describe the reversible coupling of different kinds of energy domains, such as a coupling of kinetic and potential energy domains in mechanics or an electro-magnetic coupling. Examples where one domain reversibly interacts with itself are gyroscopic forces in rigid body dynamics and self-advection of kinetic energy in fluid dynamics. The Lorentz force is another example where a kinetic energy domain interacts with itself in a way that depends on the state of a magnetic energy domain. Generalized transformers couple energy domains of the same kind. They may for instance represent mechanical levers or ideal electrical transformers. Frame transformations in multibody systems present an example where a transformation between different representations of the same energy domain is realized. Examples of kinematic constraints include joint constraints in multibody dynamics and the incompressibility constraint in fluid dynamics. Further, the parallel connection of capacitors, the series connection of inductors or springs, as well as rigid links between point masses are described by constraints.
The geometric structure behind reversible components is called Dirac structure. For our purposes, an (almost) Dirac structure $\mathcal{D}$ is a power-preserving, locally linear relation of port variables. More precisely, we could define $\mathcal{D}\subset\mathcal{P}_{I}$ as a maximally isotropic vector subbundle of the bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$, see Courant (1990); Bursztyn (2013); van der Schaft and Jeltsema (2014). However, to state fundamental properties of reversible components, we give a more practical definition of Dirac structures in terms of a particular representation, which may be called constrained hybrid input-output representation (cf. van der Schaft and Jeltsema (2014)):
@@
Let $I$ be an interface and let $\mathcal{X}_{I,p}\cong\mathcal{X}_{1}\times\mathcal{X}_{2}\ni x_{p}=(x_{1},\,x_{2})$ be a factorization of the state space of its power ports with flow and effort variables correspondingly denoted by $f=(f_{1},\,f_{2})$ and $e=(e_{1},\,e_{2})$. A Dirac structure $\mathcal{D}\subset\mathcal{P}_{I}$ admits the following representation:
@@
Here, $L:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}\mathcal{X}_{1}$, $g:\mathrm{T}\mathcal{X}_{2}\rightarrow\mathrm{T}\mathcal{X}_{1}$, its linear dual $g^{*}:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}^{*}\mathcal{X}_{2}$, $C:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathcal{X}_{I}\times V$, and its dual $C^{*}:\mathcal{X}_{I}\times V^{*}\rightarrow\mathrm{T}\mathcal{X}_{1}$ are vector bundle maps over the identity on $\mathcal{X}_{I}$. Further, $V$ is a vector space and $L(x)$ is skew-symmetric for all $x\in\mathcal{X}_{I}$. ∎
The maps $L$, $g$ and $C$ define a gyrator, a transformer and a constraint, respectively. The constraint is enforced via the Lagrange multiplier $\lambda_{c}$. Due to the skew-symmetry of the matrix in 6.7, exergetic power is conserved, i.e.
@@
Let the interface $I_{\text{hk}}=(N_{\text{hk}},\,\tau_{\text{hk}})$ be given by
$\begin{split}N_{\text{hk}}\>&=\>\{\mathtt{v_{1}},\,\mathtt{v_{2}},\,\mathtt{p}\}\\ \tau_{\text{hk}}(\mathtt{v_{1}})\>&=\>\tau_{\text{hk}}(\mathtt{v_{2}})\>=\>((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p})\\ \tau_{\text{hk}}(\mathtt{p})\>&=\>((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p})\end{split}$
and let $a\in\mathbb{R}$ be a constant surface area of a piston. Then, the following equation defines a Dirac structure $\mathcal{D}_{\text{hk}}\subset\mathcal{P}_{I_{\text{hk}}}$:
$\left[\begin{array}[]{c}\mathtt{v_{1}.f}\\ \mathtt{v_{2}.f}\\ \mathtt{p.f}\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&a\\ 0&0&-a\\ -a&a&0\end{array}\right]\,\left[\begin{array}[]{c}\mathtt{v_{1}.e}\\ \mathtt{v_{2}.e}\\ \mathtt{p.e}\end{array}\right]$
This describes a generalized gyrator that couples the two hydraulic energy domains at the front and the back side of a piston with a kinetic energy domain corresponding to the piston’s moving mass. ∎
A reversible component is defined by a Dirac structure which satisfies two extra properties ensuring thermodynamic consistency.
@@
A reversible component $(I,\,\mathcal{D})$ is given by an interface $I$ and a Dirac structure $\mathcal{D}\subset\mathcal{P}_{I}$ such that all quantities of the reference environment are conserved and the time-reversal invariance property holds. The first condition requires that
$\forall\,p_{R}\in I_{R}\>:\>\forall\,x\in\mathcal{X}_{I}\>:\>\Bigl{(}0\>=\>L(x)\,\lambda(I_{1},\,p_{R})\quad\text{and}\quad 0\>=\>g^{*}(x)\,\lambda(I_{1},\,p_{R})\Bigr{)}\,.$

### v2

@@ 5 Exergy
Every system is ultimately determined by a power-preserving interconnection of primitive subsystems, called components. These represent elementary physical behaviors, namely storage, and reversible as well as irreversible exchange of energy. To achieve a thermodynamically consistent combination of reversible and irreversible dynamics, components are defined with respect to an exergy reference environment. Before components can be defined in the next section, in this section we define the reference environment. We start with its physical interpretation.
@@ 5.1 Intuition
As briefly discussed in Section 1.4, thermodynamic models have an irreversible dynamics, since not all physically-relevant degrees of freedom are determined by the state variables. According to the first and second law of thermodynamics, internal energy, which is given by an ‘imprecise’ phenomenological description, cannot be fully converted into mechanical or electromagnetic energy, which is ‘precise’, in the sense that all relevant degrees of freedom are fully determined by the state variables.
The irreversible degradation of ‘mechanically precise’ forms of energy into ‘thermodynamically imprecise’ forms of energy can be quantified using the concept of exergy. A hypothetical conversion device, whose mode of operation is only restricted by the first and second law of thermodynamics is used to determine how much mechanical (or electromagnetic) energy can be extracted from a system, before it reaches thermodynamic equilibrium with a reference environment, see Figure 9. The environment is defined by intensive variables, such as temperature and pressure, which are constant in time and space.
@@
The ideal conversion device is not restricted to interacting with the system through its interface. Instead, the device operates in a counterfactual where all energy domains of both the system and the environment can be accessed. The exergy content of a composite system is hence obtained by (recursively) summing the exergy stored in each of its storage components. The exergy content of a component storing purely mechanical or electromagnetic forms of energy is simply equal to its energy content. For components storing internal energy, the exergy content is assessed based on the ideal conversion device and the reference environment. A concrete statement of this follows in 6.3, after the definition of the exergy storage function in 6.2.
For more information about exergy, its history, and its links to port-Hamiltonian as well as metriplectic or GENERIC systems, we refer to [3].
@@ 5.2 Reference environment
The internal energy of a thermodynamic system is a function of its entropy, and possibly further extensive quantities such as volume, mass of chemical species, etc. To assess the exergy content of the system, we need for each such variable an energy-conjugate, intensive variable attributed to the reference environment. Since these intensive variables are constant, they determine the equilibrium state of the combined system, see Figure 9.
@@
A reference environment $(I_{R},\,\lambda_{R})$ is defined by an interface $I_{R}=(N_{R},\,\tau_{R})$ whose ports are all power ports and have a unique extensive quantity in $\mathbf{Q}_{\mathbb{R}}$. One port must have the quantity $(\mathbb{R},\,\mathtt{entropy})$. Moreover, each port $p\in I_{R}$ has an associated value $\lambda_{R}(p)\in\mathbb{R}^{*}$ of the energy-conjugate intensive quantity. ∎
@@
The minimal definition of a reference environment represents an isothermal heat bath. This amounts to an interface with a single entropy port and an associated reference temperature, henceforth denoted by ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$:
$N_{R}$ =$=\>\{\mathtt{s}\}$
$\tau_{R}(\mathtt{s})$ $=\>\left((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p}\right)$
$\lambda_{R}(\mathtt{s})$ $=\>{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,.$
To determine the exergy content of pressurized fluids with respect to an isobaric atmosphere, the definition of the environment is extended to $N_{R}=\{\mathtt{s},\,\mathtt{v}\}$ with
$\tau_{R}(\mathtt{v})$ $=\>\left((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p}\right)$
$\lambda_{R}(\mathtt{v})$ $=\>{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\pi_{0}}\,.$
The minus sign in front of the reference pressure ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}}$ stems from the fact that more volume means less energy. To model for instance a mass diffusion process, a port $\mathtt{m}$ representing the mass of a chemical species with a constant chemical potential ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mu_{0}}$ can be included, i.e. let $N_{R}=\{\mathtt{s},\,\mathtt{v},\,\mathtt{m}\}$ with
@@
A software implementation should already include a comprehensive definition of a reference environment. Allowing the user to change the values of the intensive quantities would be of interest for exergy analysis and thermodynamic optimization.
@@ 6 Primitive systems
At the lowest level, any hierarchically-structured model is defined by primitive systems, called components. There are four kinds of them:
@@
    or a transformation between different forms of describing an energy domain.
@@
    Environment components may be used to model energy exchange with
@@
Components are in general defined with respect to a reference environment $(I_{R},\,\lambda_{R})$. Further, every component has an interface $I=(N,\,\tau)$ with port variables $(x,\,f,\,e)\in\mathcal{P}_{I}$. To define the exergy storage function of storage components and to state conditions that reversible and irreversible components have to satisfy, we define the following notation: For any port $p\in I$, let
@@
With regard to the existential quantifier, this is well-defined due to the injectivity of $\tau_{R}$. For instance, if $p$ represents a thermal energy domain then $\lambda(p)={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$, whereas if $p$ represents a kinetic energy domain then $\lambda(p)=0$, since the environment does not have a kinetic energy domain. We extend the above notation to an entire interface: Given for example $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
@@
We also need a similar notation that references a specific environment port: For any port $p\in I$ and $p_{R}\in I_{R}$, let
@@
Again, given for example $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
@@ 6.1 Storage components
Storage components are primitive systems that represent energy storage. Examples include elastic springs, movable masses, capacitors, inductors, and thermal capacities.
@@
The function $E$ yields the stored energy for any state $x\in\mathcal{X}_{I}$. Based on the reference environment $I_{R}$, the exergy content is determined as follows.
@@
Let $(I,\,E)$ be a storage component. Its exergy storage function $H:\mathcal{X}_{I}\rightarrow\mathbb{R}$ is defined by
@@
The exergy function $H$ follows directly from the energy function $E$ by subtracting the pairing between each state variable representing an extensive quantity also present in the environment, and the constant, energy-conjugate quantity from the environment. We come back to this in 6.3.
@@
The effort variables are given by the differential of the exergy function evaluated at the current state $x\in\mathcal{X}_{I}$, i.e.
$\forall\,p\in I\>:\>p\mathtt{.e}\>=\>\left.\partialderivative{H}{(p\mathtt{.x})}\right|_{x}\,.$
The evolution of the state variables is obtained by integrating the flow variables, i.e.
[…]
@@
Consider a storage component $(I_{\text{gas}},\,E_{\text{gas}})$ representing a gas-filled compartment whose entropy and volume can change. Its interface $I_{\text{gas}}$ is defined in 2.18, and we assume knowledge of its energy storage function $E_{\text{gas}}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$. With $x=(s,\,v)$, the exergy storage function is then given by
$H_{\text{gas}}(s,\,v)\>=\>E_{\text{gas}}(s,\,v)\,-\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,s\,+\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}}\,v\,.$
As the temperature is defined by $\theta=\partialderivative{E_{\text{gas}}}{s}$ and the (negative) pressure is defined by $-\pi=\partialderivative{E_{\text{gas}}}{v}$, the effort variables are given by
$\begin{split}\mathtt{s.e}\>&=\>\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\\ \mathtt{v.e}\>&=\>-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}})\,.\end{split}$
Since the entropy rate $\dot{s}=\mathtt{s.f}$ is related to the heat flow rate $\Phi$ via $\dot{s}=\frac{1}{\theta}\,\Phi$, the exergetic power at the entropy port is $\langle\mathtt{s.e}\mid\mathtt{s.f}\rangle=\frac{\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}{\theta}\,\Phi=\eta_{\text{Carnot}}\,\Phi$, where $\eta_{\text{Carnot}}$ is the efficiency of an ideal heat engine operating between a hot source with temperature $\theta$ and a cold sink with temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$. Considering a piston with the pressurized gas on one side and the atmosphere on the other, $\langle\mathtt{v.e}\mid\mathtt{v.f}\rangle=-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}})\,\dot{v}$ is the power that can be extracted from the pressurized gas when the volume expansion rate is $\dot{v}$. ∎
Next, we formally state the semantics of a storage component with interface $I$ in terms of a relation $\mathcal{S}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$. This is used in Section 7, where we define the semantics of a composite system in terms of the semantics of its primitive subsystems and the semantics of its (hierarchy of nested) interconnection pattern(s).
@@ 6.2 Environment components
From a user perspective, environment components provide a simple alternative to storage components, when dealing with low-grade energy.They may represent an isothermal heat bath, which absorbs heat generated by irreversible components, or an isobaric atmosphere, which exchanges pressure-volume work with other components, etc. Environment components are essentially storage components with a trivial energy and exergy function.
@@
The state variables keep track of the exchanged entropy, volume, etc. As for storage components, their rate of change is given by the flow variables, i.e.
@@
By definition, the environment has zero exergy content. Since the effort variables are given by the differential of the exergy function, they are also zero:
@@
In agreement with the definition, the chain rule also tells us that the rate of change of the stored exergy is zero.
@@
The semantics assigned to environment component $I\subseteq I_{R}$ is the relation
$\mathcal{E}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$ that is given by the subbundle
@@ 6.3 Reversible components
In a generalized sense, reversible components represent gyrators, transformers, and/or kinematic constraints. Generalized gyrators may describe a coupling of kinetic and potential energy domains in mechanics, or a coupling of magnetic and electric energy domains in electromagnetism. Examples where one domain reversibly interacts with itself are gyroscopic forces in rigid body dynamics and self-advection of kinetic energy in fluid dynamics. The Lorentz force is an example where a kinetic energy domain interacts with itself in a way that depends on the state of a magnetic energy domain. Generalized transformers may for instance describe ideal mechanical levers or electrical transformers. Frame transformations in multibody systems are an example relating different representations of the same energy domain. Examples of kinematic constraints include joint constraints in multibody dynamics and the incompressibility constraint in fluid dynamics. Further, the parallel connection of capacitors, the series connection of inductors or springs, and rigid links between point masses are described by constraints.
The geometric structure behind reversible components is called Dirac structure. For our purposes, a Dirac structure $\mathcal{D}$ is a power-preserving, locally linear relation of port variables. More precisely, we could define $\mathcal{D}\subset\mathcal{P}_{I}$ as a maximally isotropic vector subbundle of the bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$, see [1, 4, 26]. However, to state additional conditions that reversible components have to satisfy, we give a more practical definition of Dirac structures in terms of a particular representation, which might be called constrained hybrid input-output representation (cf. [4]).
@@
Let $I$ be an interface and let $\mathcal{X}_{I,p}\cong\mathcal{X}_{1}\times\mathcal{X}_{2}\ni x_{p}=(x_{1},\,x_{2})$ be a factorization of the state space of all power ports with flow and effort variables correspondingly denoted by $f=(f_{1},\,f_{2})$ and $e=(e_{1},\,e_{2})$. A Dirac structure $\mathcal{D}\subset\mathcal{P}_{I}$ admits the following representation:
@@
Here, $L:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}\mathcal{X}_{1}$, $g:\mathrm{T}\mathcal{X}_{2}\rightarrow\mathrm{T}\mathcal{X}_{1}$, its linear dual $g^{*}:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}^{*}\mathcal{X}_{2}$,
$C:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathcal{X}_{I}\times V$, and its dual $C^{*}:\mathcal{X}_{I}\times V^{*}\rightarrow\mathrm{T}\mathcal{X}_{1}$ are vector bundle maps over the identity on $\mathcal{X}_{I}$. Further, $V$ is a vector space and $L(x)$ is skew-symmetric for all $x\in\mathcal{X}_{I}$. ∎
The maps $L$, $g$ and $C$ respectively define a generalized gyrator, a generalized transformer and a constraint. The constraint is enforced via the Lagrange multiplier $\lambda_{c}$. Due to the skew-symmetry of the matrix in 6.7, exergetic power is conserved, i.e.
@@
We remark that this actually defines an almost Dirac structure, as no integrability condition, akin to the Jacobi identity for Poisson structures, is required. However, in the unconstrained or holonomically constrained case, such a condition can be checked.
@@
Let the interface $I_{\text{hkc}}=(N_{\text{hkc}},\,\tau_{\text{hkc}})$ be given by
$\begin{split}N_{\text{hkc}}\>&=\>\{\mathtt{v_{1}},\,\mathtt{v_{2}},\,\mathtt{p}\}\\ \tau_{\text{hkc}}(\mathtt{v_{1}})\>&=\>\tau_{\text{hkc}}(\mathtt{v_{2}})\>=\>((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p})\\ \tau_{\text{hkc}}(\mathtt{p})\>&=\>((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p})\end{split}$
and let $a\in\mathbb{R}$ be a constant surface area of a piston. Then, the equation
$\left[\begin{array}[]{c}\mathtt{v_{1}.f}\\ \mathtt{v_{2}.f}\\ \mathtt{p.f}\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&-a\\ 0&0&+a\\ +a&-a&0\end{array}\right]\,\left[\begin{array}[]{c}\mathtt{v_{1}.e}\\ \mathtt{v_{2}.e}\\ \mathtt{p.e}\end{array}\right]$
defines a Dirac structure $\mathcal{D}_{\text{hkc}}\subset\mathcal{P}_{I_{\text{hkc}}}$. This describes a generalized gyrator that couples the two hydraulic energy domains at the front and the back side of a piston with the kinetic energy domain of the piston’s moving mass. ∎
A reversible component is defined by a Dirac structure that satisfies two additional conditions ensuring thermodynamic consistency.
@@
A reversible component $(I,\,\mathcal{D})$ is given by an interface $I$ and a Dirac structure $\mathcal{D}\subset\mathcal{P}_{I}$ such that all quantities present in the reference environment (in particular entropy) are conserved and the time-reversal invariance property holds. The first condition requires that $\forall\,p_{R}\in I_{R}$ and $\forall\,x\in\mathcal{X}_{I}\,\colon$
$0\>=\>L(x)\,\lambda(I_{1},\,p_{R})\quad\text{and}\quad 0\>=\>g^{*}(x)\,\lambda(I_{1},\,p_{R})\quad\text{and}\quad 0\>=\>C(x)\,\lambda(I_{1},\,p_{R})\,.$